OpenStudy (anonymous):

Given f(x) = x - 5 / 3 solve for f^-1(3)

OpenStudy (anonymous):

\[f(x)=\frac{x-5}{3}\] interpret your question as "solve for \(x\):\[\frac{x-5}{3}=3\]

OpenStudy (anonymous):

Oops, I forgot to put the choices. Its: -1 14 18 21

OpenStudy (anonymous):

Okay thanks I'll try!

OpenStudy (anonymous):

just go ahead and solve it it takes exactly two steps 1) multiply by 3 2) add 5

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank youuu so mcuh :) I get it

OpenStudy (anonymous):

yw glad you do